Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Problem

Expert replies
by alsergi » Fri Aug 22, 2008 1:46 am
Please, help me with this, problem, I am a little bit confused ...


If 2 different representatives are about to be selected from a group of 10 employees and if x is the probability that both representatives are women , is x > 1/2 ?

1. More than 1 / 2 of the employees are women.
2. The probability that both representatives are men is less than 1/10.

Thanks in advance !!!
Join the discussion
Source: — Data Sufficiency |

by nikhilagrawal » Fri Aug 22, 2008 3:57 am
is the answer A?

prob will be atleat 6c2/10,
Join the discussion

Re: Probability Problem

by sudhir3127 » Fri Aug 22, 2008 4:44 am
alsergi wrote:Please, help me with this, problem, I am a little bit confused ...


If 2 different representatives are about to be selected from a group of 10 employees and if x is the probability that both representatives are women , is x > 1/2 ?

1. More than 1 / 2 of the employees are women.
2. The probability that both representatives are men is less than 1/10.

Thanks in advance !!!
I go with B

heres how...

Statement 1.
more than half are women .. Assume women to be 6 and men to be 4
selecting 2 women out of 10 will be
6c2/10c2 = 1/3 ................. Insufficient

statement 2.

probability of 2 men is < 1/10

thus probability of 2 women will be 1-(<)1/10

thus it will be >9/10..... sufficient

hence B

hope it helps...
Join the discussion

by alsergi » Fri Aug 22, 2008 5:11 am
Thank you very much sudhir3127, but I dont agree with you ...

I agree with the first statement but not with the second

1 - 1/10 would be the probability of AT LEAST one man, not the probability of 2 women so, you cant answer B ...

The right answer is E, but I dont understand exactly why ...

Thank you very much!!! I am still waiting for the right explanation, thinking it too ... :roll:
Join the discussion

by pepeprepa » Fri Aug 22, 2008 5:42 am
This trouble me during some minutes, I thought it was B first...

2)The probability that both representatives are men is less than 1/10

You fastly find that this is possible when men are either 2 or 3.
When 2 men: P=2/10*1/9=1/45
When 3 men: P=3/10*2/9=1/15
They are both less than 1/10 and when 4men the probability is above 1/10

In these cases the probability to get 2 women are:
When 2 men: W=8/10*7/9=28/45>1/2
When 3 men: W=7/10*6/9=21/45<1/2

So we can't really now, and given 2) gives us more information than 1) that's E...
Join the discussion

by rhymes_with_luck » Fri Aug 22, 2008 6:46 am
It has been discussed before.

in B commonntrait is to ignore total sample space

which can be selecting two women (or) selecting one woman and one man (or) selecting two men

As selecting one woman and one man can not be determined , hence E
Join the discussion